# Adaptive Control for Triadic Human-Robot-FES Collaboration in Gait Rehabilitation: A Pilot Study

### 01 Feb 2024 v1

@@ III-E Experimental Setup
During experimental validation of the controller, the participant was fitted with the exoskeleton and was asked to walk at their preferred speed and follow the reference path. Two 10-minute recordings were performed; one while using the HAPC and one while using the HPC with a 5-minute break between trials. Hybrid assistance was provided only for the muscles of the knee joint while for the hip joint only robotic assistance was provided (surface electrodes were placed at the quadriceps and hamstrings as shown in Fig. 1b). The position of the ankles was adjusted according to the relative position in the gait cycle. The participant’s ability to follow the reference path, the assistance they received and their estimated muscle fatigue were quantified and the sum of these parameters was calculated in order to evaluate the controller’s ability to provide assistance as needed and prevent muscle fatigue. To facilitate the comparison between controllers, the results were normalised such that the costs for the HAPC are equal to 1.
@@ IV-A Simulation Results
Fig. 3 shows the controller’s response to the simulated human behaviour described in section III-D. It can be seen that for the first half of the simulation (indicated by a grey dotted line) the adaptive controller results in a reduction in exoskeleton stiffness, exoskeleton assistance (Fig. 3a), and FES stiffness through the adaptation of parameter $\gamma$ (Fig. 3b), while the trajectory error does not significantly change (Fig. 3b) and muscle fatigue slowly increases (Fig. 3c). For the second half of the simulation the improved performance of the model reduces the tracking error (Fig. 3b) and the assistance received from both the exoskeleton (Fig. 3a) and the FES (Fig. 3b-c), which allows the muscles to recover (Fig. 3c).
In Fig. 4 the performance of four different controllers is compared based on their ability to reduce the trajectory error, the assistance provided by the exoskeleton and the FES, the mean of muscle fatigue, and the sum of the four. It can be seen that, compared to the HAPC, the EPC provides almost six times as much robotic assistance to keep the trajectory error lower while the FPC results in higher trajectory error, higher stimulation intensity and more muscle fatigue than the HAPC, resulting in a higher sum of the normalised costs. When comparing the HAPC to the HPC, a visible reduction in robotic assistance is observed while the rest of the parameters are affected less. This indicates that the hybridisation of robotic assistance and FES in this hierarchical order, and the controller’s adaptation, provide a substantial reduction in the assistance provided to the user (at the expense of a slightly higher tracking error when compared to the EPC), which overall reduces the sum of error, assistance and muscle fatigue, and implies an improved ability to provide assistance as needed.
@@ IV-B Experimental Results
Fig. 5 shows the response of the HAPC to the recorded human behaviour and Fig. 6 provides a comparison between the HAPC and the HPC. Based on the results presented in Fig. 5a it is evident that the controller’s adaptation resulted in a significantly lower exoskeleton stiffness than the baseline stiffness for both the hips and the knees [^1] Results are shown for the left leg only as the results for the right leg follow similar patterns. These results can be provided upon requests.. This decrease in exoskeleton stiffness was also reflected in the assistance provided by the robot (Fig. 5a). It is also evident that the exoskeleton stiffness of the knee joint converged to a smaller value than the exoskeleton stiffness of the hip joint (Fig. 5a), which could be indicative of the cooperation between the robot and the FES. Similarly, a reduction in parameter $\gamma$ is observed, which reduces the stiffness of the FES controller (Fig. 5b). This is also reflected in the stimulation provided by the adaptive controller to the knee (Fig. 5c). As a result, a decreased muscle fatigue is observed (Fig. 6), which for the left leg was more pronounced for the knee extensor muscles (Fig. 5c). Meanwhile, the trajectory tracking error of the knee joint seems to vary only slightly (Fig. 5b).

### 08 Mar 2024 v2

@@ III-E Experimental Setup
During experimental validation of the controller, the participant was fitted with the exoskeleton and was asked to walk at their preferred speed and follow the reference path as accurately as possible. Two 10-minute recordings were performed with the aid of visual feedback; one while using the HAPC and one while using the HPC with a 5-minute break between trials. Hybrid assistance was provided only for the muscles of the knee joint while for the hip joint only robotic assistance was provided (surface electrodes were placed at the quadriceps and hamstrings as shown in Fig. 1b). The position of the ankles was adjusted according to the relative position in the gait cycle. The participant’s ability to follow the reference path, the assistance they received and their estimated muscle fatigue were quantified and the sum of these parameters was calculated in order to evaluate the controller’s ability to provide assistance as needed and prevent muscle fatigue. To facilitate the comparison between controllers, the results were normalised such that the costs for the HAPC are equal to 1.
@@ IV-A Simulation Results
Fig. 3 shows the controller’s response to the simulated human behaviour described in section III-D. It can be seen that for the first half of the simulation (indicated by a grey dotted line) the adaptive controller results in a reduction in exoskeleton stiffness, exoskeleton assistance (Fig. 3a), and FES stiffness through the adaptation of parameter $\gamma$ (Fig. 3b), while the trajectory error follows a periodic pattern (Fig. 3b) and muscle fatigue slowly increases (Fig. 3c). For the second half of the simulation the improved performance of the model reduces the tracking error (Fig. 3b) and the assistance received from both the exoskeleton (Fig. 3a) and the FES (Fig. 3b-c), which allows the muscles to recover (Fig. 3c).
In Fig. 4 the performance of four different controllers is compared in response to the simulated human behaviour described in section III-D. It can be seen that, compared to the HAPC, the EPC provides almost six times as much robotic assistance to keep the trajectory error low while the FPC results in higher trajectory error, higher stimulation intensity and more muscle fatigue than the HAPC. When comparing the HAPC to the HPC, a visible reduction in robotic assistance is observed while the rest of the parameters are affected less. This suggests that sufficient assistance is provided by the FES and that the controller’s adaptation can reduce excessive assistive forces from the exoskeleton. These observations indicate that the hybridisation of robotic assistance and FES in this hierarchical order, and the controller’s adaptation, provide a substantial reduction in the assistance provided to the user (at the expense of a higher tracking error when compared to the EPC), which overall reduces the sum of error, assistance and muscle fatigue, and implies an improved ability to provide assistance as needed.
@@ IV-B Experimental Results
Fig. 5 shows the response of the HAPC to the recorded human behaviour and Fig. 6 provides a comparison between the HAPC and the HPC. Based on the results presented in Fig. 5a it is evident that the controller’s adaptation converged towards a lower exoskeleton stiffness than the baseline stiffness for both the hips and the knees [^1] Results are shown for the left leg only as the results for the right leg follow similar patterns. These results can be provided upon request.. This decrease in exoskeleton stiffness was also reflected in the assistance provided by the robot (Fig. 5a). It is also evident that the exoskeleton stiffness of the knee joint converged to a smaller value than the exoskeleton stiffness of the hip joint (Fig. 5a), which could be indicative of the cooperation between the robot and the FES. Similarly, a reduction in parameter $\gamma$ is observed, which reduces the stiffness of the FES controller (Fig. 5b). This is also reflected in the stimulation provided by the adaptive controller to the knee (Fig. 5c). As a result, a decreased muscle fatigue is observed (Fig. 6), which for the left leg was more pronounced for the knee extensor muscles (Fig. 5c). Meanwhile, the trajectory tracking error of the knee joint seems to vary only slightly (Fig. 5b).